OpenStudy (anonymous):

Why \(lim_{n \to \infty}(1-\frac{1}{n})^{n} = e^{-1} \) ? OR How can i find this rule, whats it name ?

OpenStudy (anonymous):

It is very well known that \[ \lim_{n->\infty} \left ( (1+\frac x n \right)^n = e^x \] put x=-1 and you are done.

OpenStudy (anonymous):

Let \[ a_n= \left ( 1+\frac x n \right)^n\\ \ln(a_n) = n \ln \left ( (1+\frac x n \right)\\ \ln(a_n) = \frac{ \ln \left ( (1+\frac x n \right)}{\frac 1 n}\\ \] Apply L'Hospital's Rule (Derivative with respect to n) you get\[ \lim_{n->\infty} \ln a_n= \lim_{n->\infty} -\frac{x}{\frac{(-1) \left(n^2 \left(\frac{x}{n}+1\right)\right)}{n ^2}}=\\ \lim_{n->\infty} \frac{x}{\frac{x}{n}+1}=x\\ \lim_{n->\infty} a_n =e^x \]